OpenStudy (secret-ninja):

STOP CLOSING MY QUESTION IM GETTING HELP The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3 times f(n -1) - 2 times f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, ... -4, -12, -28, -60, ... -4, 12, -18, 54, ... -4, 12, -18, -54, ...

OpenStudy (secret-ninja):

I know its not B, because it is 12, not -12.

OpenStudy (amistre64):

i spose .... it looked spammy at first

OpenStudy (amistre64):

yeah, B is out

OpenStudy (amistre64):

so the third term will settle it for us right?

OpenStudy (secret-ninja):

I'm sorry, I only kept reposting because the mod kept closing it, and someone was finally helping me and they deleted it! Its so frustrating, that mod didn't want me to get help!

OpenStudy (secret-ninja):

You mean option 3?

OpenStudy (amistre64):

no, the third term is f(2) so let n=2 and work the rule

OpenStudy (amistre64):

if its -18, then we need to determine the last term of f(3)

OpenStudy (secret-ninja):

Is it D?

OpenStudy (amistre64):

you need to work the math, afterwards the solution will be obvious

OpenStudy (amistre64):

f(n) = -3 times f(n -1) - 2 times f(n - 2) f(2) = -3 times f(2-1) - 2 times f(2-2) = ???

OpenStudy (secret-ninja):

-18?

OpenStudy (amistre64):

most likely f(2) = -3 f(1) -2 f(2) = -3(-4) -2(12) = 12 - 24 = -12 so im most likely not reading your posted 'format' correctly

OpenStudy (amistre64):

can you fix your math notation?

OpenStudy (amistre64):

and of course that f(0) on the end up there, typo

OpenStudy (secret-ninja):

I don't understand, are you saying I may have typed something wrong in my question?

OpenStudy (secret-ninja):

I didn't type anything differently from my real question :/

OpenStudy (amistre64):

yes. i get -12 as the 3rd term so either your options are inccorect, or i have misread the rule for f(n) since its not in a proper math notation format

OpenStudy (amistre64):

f(n) = -3 times f(n -1) - 2 times f(n - 2) this is confusing to try to read properly as a math equation

OpenStudy (secret-ninja):

how about this: \[f(n)=-3 x f(n-1) -2 x f(n-2)\]

OpenStudy (secret-ninja):

the x aren't variables, just multiplication signs

OpenStudy (amistre64):

let me retry this .. it could be im readingit rightand just placing values in the wrong spot f(0) = -4 f(1) = 12 f(n) = -3 times f(n -1) - 2 times f(n - 2) f(2) = -3 times f(2 -1) - 2 times f(2 - 2) f(2) = -3 times f(1) - 2 times f(0) f(2) = -3(12) - 2(-4) -36 + 8 = -28

OpenStudy (secret-ninja):

So it would be A or B

OpenStudy (amistre64):

thats better, so we want the 3rd tern to be -28 which narrows it to one options since B is bad

OpenStudy (secret-ninja):

So its A?

OpenStudy (amistre64):

yes, id go with A

OpenStudy (secret-ninja):

Thank you so much! :D

OpenStudy (amistre64):

yep